OBSERVATION OF AN UNGERADE RYDBERG STATE OF ^{7}Li_{2}

\underline{A. J. ROSS}, F. MARTIN and A. ADOHI-KROU^{a}, Laboratoire de Spectrométrie Ionique et Moléculaire (UMR 5579 CNRS), Bâtiment A. Kastler, Université Lyon I, Domaine Scientifique de la Doua, 69622 Villeurbanne Cedex, France.

Some weak, collisionally induced transitions have been observed following excitation of the two lowest vibrational levels of the 5d ^{1}Π_{g} state of ^{7}Li_{2} (produced in a heatpipe source). The transitions were recorded in Fourier transform resolved fluorescence spectra in the near infrared (7900 - 8800 cm^{−1}), at an instrumental resolution of 0.08 cm^{−1}, and they have been assigned as transitions from a Rydberg ^{1}Π_{u} state to v = 0, 1 and 2 of the 1 ^{1}Δ_{g} electronic state. It seems highly likely that the two upper state levels populated through the collisions are v = 0 and v = 1 of a ^{1}Π_{u} state which lies very close in energy to the 5d ^{1}Π_{g} state. The energy transfer occurs to v = 0 of the ^{1}Π_{u} state from v = 0 of the 5d ^{1}Π_{g} state, and (at low J only) to v = 1 of the ^{1}Π_{u} state from v = 1 of 5d ^{1}Π_{g}. Six bands were analysed: 0-0, 0-1, 0-2, 1-0, 1-1 and 1-2. The rotational constants of this ^{1}Π_{u} state are too large to be high levels of the other ungerade Rydberg state of Li_{2} documented in the literature, namely the D ^{1}Π_{u} state studied by Theiss and co-workers^{b}. They are however sufficiently close to those of the 5d ^{1}Π_{g} state (T_{e} = 37257.7 cm^{−1}, B_{e} = 0.48 cm^{−1}, ω_{e} = 237.4 cm^{−1}) given in earlier work^{c}, that we suggest the Rydberg state is most likely to correlate at long range with Li(2s) + Li(5p) atoms. Principal spectroscopic constants are T_{0} = 37390.198(5) cm^{−1}, B_{0} = 0.50627(12) cm^{−1}, ΔG_{v} = 262.4 cm^{−1}.

aPermanent address: UFR SSMT Université de Cocody, 22 BP 582 Abidjan 22, Ivory Coast.

bW. Theiss, H. J. Müschenborn, and W. Demtröder, Chem. Phys. Lett. \underline{174} (1990) 126-132

cR. Bernheim, L. P. Gold, and T. Tipton, J. Chem. Phys. \underline{78} (1983) 3635-3646